OpenStudy (anonymous):

What is the equation of the line perpendicular to the line y=4/9x-2 that passes through the point (4,3)?

OpenStudy (anonymous):

To find perpendicular lines use this : m1.m2=-1 Means for example if we have a line : y=8x+3 m2=-1/8 OK ?:)

OpenStudy (anonymous):

@mrivera3113 :Got it ????

OpenStudy (anonymous):

I still don't get it?

OpenStudy (anonymous):

amm.. To have two line perpendicular if we have m(for first line ) * m(for 2th line ) = -1 Got it ?:)

OpenStudy (anonymous):

So would it be y= -9 / 4 x+12?

OpenStudy (anonymous):

Is it an answer for all of the question ????

OpenStudy (anonymous):

the answer choices for the question is y= -4 / 9 x - 2 y = 9 / 4 x + 6 y= -9 / 4 x + 12 y= -9 / 4 x - 12

OpenStudy (anonymous):

@E.ali

OpenStudy (anonymous):

OK ! I asked u another ! No problem ! Did u get to how find slope(m) ?

OpenStudy (anonymous):

@mrivera3113

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Look : If we have this equation answer : 9 . x = -1 x= ?

OpenStudy (anonymous):

@mrivera3113

OpenStudy (anonymous):

slope of th required line will be -9/4 bcoz m1*m2=-1

OpenStudy (anonymous):

the required eqn is, (y-y1)=m(x-x1) (y-3)=(-9/4)(x-4) 4(y-3)=-9(x-4) 4y-12=-9x+36 4y+9x=36+12 4y+9x=48

OpenStudy (anonymous):

So it would be -9 / 4 x + 12? Sorry it took forever to reply I was at work.

OpenStudy (anonymous):

yes the third option is the correct answer